### v2

@@ Abstract
Inherent temporal heterogeneity, such as varying sampling densities and periodic structures, has posed substantial challenges in zero-shot generalization for Time Series Foundation Models (TSFMs). Existing TSFMs predominantly rely on massive parameterization to absorb such heterogeneity, as their static tokenization and positional encoding schemes entangle diverse temporal patterns into a fixed representation space, encouraging memorization rather than adaptation. To address this limitation, we propose Kairos, a flexible and parameter-efficient TSFM that decouples temporal heterogeneity from model capacity through a novel tokenization perspective. Kairos introduces a dynamic patching tokenizer and a mixture-of-size encoding that adapt observational granularity to local information density, enabling fine-grained temporal abstraction without increasing model width or depth. In addition, we design a multi-granularity positional embedding based on dynamic rotary encodings, which conditions on instance-level spectral features and temporal structure induced by dynamic patching tokenization, allowing robust modeling of diverse temporal dependencies. Trained on a novel Predictability-Stratified Time-Series (PreSTS) corpus, Kairos achieves superior zero-shot performance with substantially fewer parameters on two mainstream benchmarks, GIFT-Eval and Time-Series-Library. The project page is at https://foundation-model-research.github.io/Kairos.
Figure 1: (a) Comparison on the GIFT-Eval benchmark (Aksu et al., 2024) demonstrates that our Kairos achieves superior zero-shot forecasting performance (lower normalized MASE) while requiring significantly fewer parameters than existing TSFMs. (b) (c) Significant variation exists in information density (i.e., signal complexity) across and within different time series datasets. (d) Existing TSFMs primarily use tokenization methods like point-wise or fixed-size patching, while our Kairos uses Mixture-of-Size Tokenization to dynamically adapt to information density, enabling the learning of generalizable rules rather than memorization.
@@ 1 Introduction
Time series forecasting is a core component of many real-world applications (Box et al., 2015; Wang et al., 2024b). Although time series data are ubiquitous, many individual datasets suffer from scarcity, making it difficult to train models independently, particularly for deep neural networks that require large amounts of data. This challenge has long motivated the development of few-shot and zero-shot forecasting methods (Oreshkin et al., 2020). More recently, inspired by the success of large-scale foundation models (Achiam et al., 2023; Kirillov et al., 2023), researchers have turned to building Time Series Foundation Models (TSFMs) for forecasting. These models leverage vast and diverse time series corpora to acquire strong generalization capabilities, showing promising results across various downstream tasks (Ansari et al., 2024; Woo et al., 2024; Shi et al., 2024).
@@
Despite this progress, a key limitation persists: current TSFMs rely on non-adaptive processing pipelines that fail to reflect the dynamic and heterogeneous nature of time series, severely compromising their parameter efficiency. As shown in Figure 1(b)(c), the statistical information density (details in Appendix G) varies not only across datasets but also across segments within the same series. By applying rigid architectures (exemplified in Figure 1(d)) to this inherently heterogeneous data, models are forced to use massive parameters to memorize specific patterns rather than acquiring adaptive, generalizable rules. Unfortunately, this strategy diminishes the overall utility of the model’s capacity, rendering it less effective for real-world deployment.
@@
This deficiency manifests primarily in two fundamental aspects of model design: (i) Representation learning: TSFMs have yet to adopt dynamic tokenization strategies analogous to those in Large Language Models (LLMs). While LLMs utilize subword tokenization (Gage, 1994) to efficiently compress semantic units of varying lengths, TSFMs predominantly persist with static patching, whether through point-wise tokenization (Ansari et al., 2024; Shi et al., 2024) or fixed-size patches (Das et al., 2024). By coupling token generation to fixed time intervals rather than information content, these approaches forcibly compress diverse temporal variations into a rigid representation space. Consequently, the model wastes computational capacity on low-information segments while struggling to capture high-frequency details within the same fixed budget, leading to significant parameter inefficiency. (ii) Temporal dependency modeling: Conventional positional encodings in TSFMs impose a unified temporal scale across sequences, neglecting differences in periodicity, trend and seasonality. As a result, existing TSFMs like (Liu et al., 2025) struggle to adapt to heterogeneous domains, for example, power consumption data measured hourly versus retail sales data measured daily, each exhibiting distinct temporal dynamics, as we empirically validate in Section 4.3. Consequently, the model requires a deeper, more complex architecture to compensate for this misalignment, further exacerbating parameter inefficiency and hindering practical utility.
@@
To address these limitations, we propose Kairos, a parameter-efficient model comprising a Mixture-of-Size Encoder, a Heterogeneity-Aware Transformer, and a Multi-Patch Decoder. Kairos addresses time series heterogeneity through two primary mechanisms: (i) For intra-patch representation learning, the Mixture-of-Size Encoder adaptively models time series at multiple granularities based on local characteristics. Inspired by null experts in Mixture-of-Experts (Zeng et al., 2024; Jin et al., 2024), we introduce null patch sizes to dynamically adjust the number of active granularities. This approach moves beyond static tokenization strategies, enabling more expressive modeling of nuanced, time-varying dynamics. (ii) For inter-patch dependency modeling, we enhance the Transformer with Dynamic Rotary Position Embedding (DRoPE). Unlike standard RoPE (Su et al., 2024), which represents positional information with a fixed temporal scale, DRoPE modulates temporal scales using spectral features and granularity to generate instance-specific positional encodings. Flexible intra-patch and inter-patch modeling enable Kairos to effectively characterize heterogeneous time series with a substantially reduced parameter count (see Figure 1(a)). Additionally, the Multi-Patch Decoder uses learnable tokens to predict future patches in parallel, which mitigates cumulative errors in autoregressive generation and offers greater flexibility for variable-length prediction horizons.
@@
To further enhance training, we construct the Predictability-Stratified Time Series (PreSTS) corpus, a large-scale and diverse pretraining dataset curated through a targeted sampling strategy. By prioritizing more predictable sequences while maintaining broad coverage, PreSTS provides high-quality supervision that supports efficient model scaling. As demonstrated by the results on the GIFT-Eval benchmark (Aksu et al., 2024) in Figure 1(a), these combined innovations allow our Kairos to achieve superior performance while using fewer parameters. This strong performance is consistently observed on Time-Series-Library (TSLib) (Wang et al., 2024b) benchmarks as well.
@@
In summary, our main contributions are as follows: (i) We present Kairos, a parameter-efficient time series foundation model designed to handle the dynamic and heterogeneous nature of time series data. (ii) We introduce three key architectural innovations: the Mixture-of-Size Encoder for adaptive multi-granularity tokenization, DRoPE for instance-specific temporal calibration, and the Multi-Patch Decoder for flexible and efficient parallel forecasting. (iii) Extensive empirical evaluations on the GIFT-Eval and Time-Series-Library benchmarks confirm the efficacy of Kairos, highlighting its superior zero-shot forecasting performance.
@@ 2 Related Work
Time Series Foundation Models (TSFMs). Inspired by the strong generalization of Large Language Models (LLMs), researchers have devoted increasing attention to TSFMs. Existing approaches typically involve adapting LLMs to time series (Gruver et al., 2023), training scalable Transformers on large corpora (Ansari et al., 2024; Das et al., 2024; Woo et al., 2024; Cohen et al., 2025; Liu et al., 2025), or exploring alternative architectures, including non-Transformer models (Auer et al., 2025; Graf et al., 2025) and lightweight Multi-Layer Perceptrons (Ekambaram et al., 2024). However, these models often apply fixed tokenization strategies and position embedding to time series from diverse domains with distinct characteristics, limiting their cross-domain generalization. Kairos, in contrast, employs a dynamic patching tokenizer and instance-adaptive, granularity-aware position embedding to enhance its generalizability for diverse time series data.
@@
Time Series Tokenization. Transformer-based models require converting raw data into discrete or continuous tokens (Sennrich et al., 2016; Agarwal et al., 2025). Current strategies generally fall into two static paradigms: (i) Fixed-size: The dominant paradigm relies on a uniform patch size across the sequence. This ranges from fine-grained point-wise tokenization (Ansari et al., 2024; Rasul et al., 2023; Shi et al., 2024) to coarser uni-size patching (Nie et al., 2023), which has been adopted by recent foundation models (Woo et al., 2024; Liu et al., 2025). While the latter improves upon the computational inefficiency of point-wise methods, both remain rigid. (ii) Multi-size: To capture broader temporal dynamics, models like Pathformer (Chen et al., 2024) and TTM (Ekambaram et al., 2024) incorporate multiple predefined patch sizes. However, these scales are applied statically, rather than adapting to the time series instance. All the above approaches partition sequences based on fixed time intervals rather than local information. This rigidity forces the model to allocate equal parameters to segments of varying complexity. This imbalance inevitably leads to suboptimal parameter efficiency and limited generalization. In contrast, Kairos employs a mixture-of-size strategy. By allocating fewer tokens to simple patterns and more to complex regions, Kairos significantly enhances parameter efficiency while maintaining robust generalization.
@@
Position Embedding. Due to the position-unaware characteristic of self-attention mechanism, Transformers Vaswani et al. (2017) rely on position embeddings (PEs) to model temporal information. However, most TSFMs adopt existing designs for PE in Natural Language Processing (NLP) domains (Das et al., 2024; Shi et al., 2024), which typically emphasize long-term decay through methods like sinusoidal calculations or by suppressing long-range positional information (Vaswani et al., 2017; Su et al., 2024). By imposing a uniform temporal scale, these PEs struggle to model the heterogeneous temporal relationships of time series. Existing adaptive PE methods in NLP (Zheng et al., 2024; Lin et al., 2024) are also insufficient, as they target context extrapolation rather than the complex temporal structures of time series. Therefore, ElasTST (Zhang et al., 2024) proposed a tunable RoPE to better adapt to time series data. However, its adaptation uses per-dataset training rather than dynamic adjustment. To address this limitation, we propose to dynamically modulate PE tailored to the intrinsic characteristics of each input time series, enhancing the parameter efficiency and overall performance of TSFMs.
@@ 3.1 Overall Framework
We formulate our task as learning a time series forecasting function $f_{\phi}:\mathbb{R}^{T}\to\mathbb{R}^{H}$ parameterized by $\phi$. Given historical time series $\mathbf{X}\in\mathbb{R}^{T}$, time series forecasting is to predict the future horizon $\mathbf{Y}\in\mathbb{R}^{H}$. The learning objective is to minimize the discrepancy between the ground truth $\mathbf{Y}$ and the prediction $\hat{\mathbf{Y}}=f_{\phi}(\mathbf{X})$. We detail it in Appendix D.1.2.
@@ 3.2 Mixture-of-Size Encoder
Granularity. For a given time series segment $\mathbf{x}$, we adopt a multi-granularity modeling perspective. Granularity refers to the temporal resolution at which the segment is tokenized and encoded, as defined below.
@@
Step 1: Top-$K$ Granularity Routing. Applying encoding at an appropriate granularity level for each time series segment is non-trivial due to the inherent temporal heterogeneity of time series data. To address this challenge, we introduce a dynamic routing mechanism equipped with a lightweight router that adaptively selects an optimal subset of granularity levels for subsequent encoding of a given segment. Inspired by the Mixture-of-Experts (MoE) paradigm (Jacobs et al., 1991; Shazeer et al., 2017), we instantiate $G$ granularity-specific experts, each corresponding to a distinct granularity level, together with $Z\in\{1,\dots,K-1\}$ null experts that perform no computation. The inclusion of null experts allows the model to implicitly skip encoding at certain granularities, enabling adaptive control over the effective number of active granularities for each input segment. The routing weights are computed via a learnable scoring function. Specifically, for each expert $i\in\{1,\dots,G+Z\}$, an affinity score is first computed as $s_{i}=\mathbf{w}_{i}^{\top}\mathbf{x},$ which is then normalized with a bias-corrected softmax $\tilde{s}_{i}=\left(\exp(s_{i}+\boldsymbol{b}_{i})\right)/\left(\sum_{j=1}^{G+Z}\exp(s_{j}+\boldsymbol{b}_{j})\right),$ where $\mathbf{w}_{i}$ denotes learnable parameters and $\boldsymbol{b}_{i}$ is a bias term introduced for auxiliary-loss-free load balancing (Liu et al., 2024a) (details in Appendix D.3.1). The router thus produces a set of active granularity indices $\mathcal{I}_{\mathrm{valid}}=\operatorname*{argtop\text{-}K}_{i\in[G+Z]}\penalty 10000\ \tilde{s}_{i}\cap\{1,\dots,G\}.$ That is, only the Top-$K$ granularity experts are activated, while null experts contribute no computation even when selected. By selectively activating granularity experts and allowing null experts to absorb routing probability mass, Kairos dynamically adapts the effective granularity resolution for each time series segment according to its local information, while maintaining a sparse and parameter-efficient architecture.
@@
where $\pi_{i}$ denotes a granularity-specific projection function, implemented as a two-layer multilayer perceptron (MLP).
Step 3: Mixture-of-Size Fusion. In the final encoder stage, we fuse the embeddings produced by the activated granularity experts into a unified token sequence, serving as the input to the Heterogeneity-Aware Transformer. First, for a time-series segment $\mathbf{x}$, we align all granularity-specific embeddings $\boldsymbol{E}_{i}$ to the finest activated resolution $P_{\min}=\min\{P_{i}\mid i\in\mathcal{I}_{\mathrm{valid}}\},$ yielding a target sequence length $M=\lceil P/P_{\min}\rceil$. This alignment is crucial to eliminate representation bias arising from unequal sequence lengths induced by different patch sizes across granularities. To enable efficient alignment, we design the patch sizes to be nested, i.e., $P_{i+1}$ is divisible by $P_{i}$. Under this constraint, each embedding sequence $\boldsymbol{E}_{i}\in\mathbb{R}^{M_{i}\times D_{h}}$ can be deterministically expanded to the target resolution via repetition (details and an example in Appendix D.3.2): $\tilde{\boldsymbol{E}}_{i}=\mathrm{Expand}(\boldsymbol{E}_{i})\in\mathbb{R}^{M\times D_{h}}.$ The fused representation is obtained by a gated weighted summation over the activated granularities: $\bar{\boldsymbol{E}}=\sum_{i\in\mathcal{I}_{\mathrm{valid}}}\alpha_{i}\,\tilde{\boldsymbol{E}}_{i},\alpha_{i}=\tilde{s}_{i}/{\sum_{j\in\mathcal{I}_{\mathrm{valid}}}\tilde{s}_{j}}.$
@@
Unlike prior approaches such as Chen et al. (2024); Zhang et al. (2024), which impose a static multi-scale partitioning over the entire time series, our method (i) dynamically adapts the tokenization granularity at the segment level. (ii) decouples temporal heterogeneity from a fixed representational scale by explicitly and dynamically assigning patch size via routing and multi-granularity encoding, enabling more effective intra-patch representation learning. This adaptive behavior cannot be achieved via globally predefined scales applied statically and is empirically validated in Section 4.4.1. As a result, varying temporal patterns are modeled via the lightweight Mixture-of-Size Encoder, reducing the burden of subsequent model parameterization.

### v3

@@ Abstract
Inherent temporal heterogeneity, such as varying sampling densities and periodic structures, has posed substantial challenges in zero-shot generalization for Time Series Foundation Models (TSFMs). Existing TSFMs predominantly rely on massive parameterization to absorb such heterogeneity, as their static tokenization and positional encoding schemes entangle diverse temporal patterns into a fixed representation space, encouraging memorization rather than adaptation. To address this limitation, we propose Kairos, a flexible and parameter-efficient TSFM dedicated to forecasting tasks, which decouples temporal heterogeneity from model capacity through a novel tokenization perspective. Kairos introduces a dynamic patching tokenizer and a mixture-of-size encoding that adapt observational granularity to local information density, enabling fine-grained temporal abstraction without increasing model width or depth. In addition, we design a multi-granularity positional embedding based on dynamic rotary encodings, which conditions on instance-level spectral features and temporal structure induced by dynamic patching tokenization, allowing robust modeling of diverse temporal dependencies. Trained on a novel Predictability-Stratified Time-Series (PreSTS) corpus, Kairos achieves superior zero-shot performance with substantially fewer parameters on two mainstream benchmarks, GIFT-Eval and Time-Series-Library. The project page is at https://foundation-model-research.github.io/Kairos.
Figure 1: (a) Comparison on the GIFT-Eval benchmark [3] demonstrates that our Kairos achieves superior zero-shot forecasting performance (lower normalized MASE) while requiring significantly fewer parameters than existing TSFMs. (b) (c) Significant variation exists in information density (i.e., signal complexity) across and within different time series datasets. (d) Existing TSFMs primarily use tokenization methods like point-wise or fixed-size patching, while our Kairos uses Mixture-of-Size Tokenization to dynamically adapt to information density, enabling the learning of generalizable rules rather than memorization.
@@ 1 Introduction
Time series forecasting is a core component of many real-world applications [6, 39]. Although time series data are ubiquitous, many individual datasets suffer from scarcity, making it difficult to train models independently, particularly for deep neural networks that require large amounts of data. This challenge has long motivated the development of few-shot and zero-shot forecasting methods [30]. More recently, inspired by the success of large-scale foundation models [1, 21], researchers have turned to building Time Series Foundation Models (TSFMs) for forecasting. These models leverage vast and diverse time series corpora to acquire strong generalization capabilities, showing promising results across various downstream tasks [4, 40, 34].
@@
Despite this progress, a key limitation persists: current TSFMs for forecasting rely on non-adaptive processing pipelines that fail to reflect the dynamic and heterogeneous nature of time series, severely compromising their parameter efficiency. As shown in Figure 1(b)(c), the statistical information density (details in Appendix K) varies not only across datasets but also across segments within the same series. By applying rigid architectures (exemplified in Figure 1(d)) to this inherently heterogeneous data, models are forced to use massive parameters to memorize specific patterns rather than acquiring adaptive, generalizable rules. Unfortunately, this strategy diminishes the overall utility of the model’s capacity, rendering it less effective for real-world forecasting applications.
@@
This deficiency manifests primarily in two fundamental aspects of model design: (i) Representation learning: TSFMs have yet to adopt dynamic tokenization strategies analogous to those in Large Language Models (LLMs). While LLMs utilize subword tokenization [14] to efficiently compress semantic units of varying lengths, TSFMs predominantly persist with static patching, whether through point-wise tokenization [4, 34] or fixed-size patches [11]. By coupling token generation to fixed time intervals rather than information content, these approaches compress diverse temporal variations into a rigid representation space. Consequently, the model wastes computational capacity on low-information segments while struggling to capture high-frequency details within the same fixed budget, leading to significant parameter inefficiency. (ii) Temporal dependency modeling: Conventional positional encodings in TSFMs impose a unified temporal scale across sequences, neglecting differences in periodicity, trend and seasonality. As a result, existing TSFMs struggle to adapt to heterogeneous domains, for example, power consumption data measured hourly versus retail sales data measured daily, each exhibiting distinct temporal dynamics, as we validate in Section 4.3. Consequently, the model requires a deeper, more complex architecture to compensate for this misalignment, further exacerbating parameter inefficiency and hindering practical utility.
@@
To address these limitations, we propose Kairos, a parameter-efficient model comprising a Mixture-of-Size Encoder, a Heterogeneity-Aware Transformer, and a Multi-Patch Decoder. Kairos addresses time series heterogeneity through two primary mechanisms: (i) For intra-patch representation learning, the Mixture-of-Size Encoder adaptively models time series at multiple granularities based on local characteristics. Inspired by null experts in Mixture-of-Experts [43, 20], we introduce computation-free null experts to dynamically adjust the number of active granularities. This approach moves beyond static tokenization strategies, enabling more expressive modeling of nuanced, time-varying dynamics. (ii) For inter-patch dependency modeling, we enhance the Transformer with Dynamic Rotary Position Embedding (DRoPE). Unlike standard RoPE [35], which represents positional information with a fixed temporal scale, DRoPE modulates temporal scales using spectral features and granularity to generate instance-specific positional encodings. Flexible intra-patch and inter-patch modeling enable Kairos to effectively characterize heterogeneous time series with a substantially reduced parameter count (see Figure 1(a)). Additionally, the Multi-Patch Decoder uses learnable tokens to predict future patches in parallel, which mitigates cumulative errors in autoregressive generation and offers greater flexibility for variable-length prediction horizons.
@@
To complement our architecture with high-quality supervision, we construct the Predictability-Stratified Time Series (PreSTS) corpus, a large-scale and diverse pretraining dataset curated through a targeted sampling strategy. By prioritizing more predictable sequences while maintaining broad coverage, PreSTS provides high-quality supervision that supports efficient model scaling. As demonstrated by the results on the GIFT-Eval benchmark [3] in Figure 1(a), these combined innovations allow our Kairos to achieve superior performance while using fewer parameters. This strong performance is consistently observed on Time-Series-Library (TSLib) [39] benchmarks as well.
@@
In summary, our main contributions are as follows: (i) We present Kairos, a parameter-efficient time series foundation model that introduces a novel architectural framework to handle the dynamic and heterogeneous nature of time series data. (ii) We curate the PreSTS corpus, a large-scale pre-training dataset of over 300 billion time points, featuring a predictability-based sampling strategy to provide high-quality and diverse supervision. (iii) Extensive empirical evaluations on the GIFT-Eval and Time-Series-Library benchmarks confirm the efficacy of Kairos, highlighting its superior zero-shot forecasting performance.
@@ 2 Related Work
Time Series Foundation Models (TSFMs) for Forecasting. Inspired by the strong generalization of Large Language Models (LLMs), researchers have devoted increasing attention to TSFMs. Existing approaches typically involve adapting LLMs to time series [18], training scalable Transformers on large corpora [4, 11, 40, 10, 28], or exploring alternative architectures, including non-Transformer models [5, 17] and lightweight Multi-Layer Perceptrons [13]. However, these models often apply fixed tokenization strategies and position embedding to time series from diverse domains with distinct characteristics, limiting their cross-domain generalization. Kairos, in contrast, employs a dynamic patching tokenizer and instance-adaptive, granularity-aware position embedding to enhance its generalizability for diverse time series data.
@@
Time Series Tokenization. Transformer-based models require converting raw data into discrete or continuous tokens [32, 2]. Current strategies generally fall into two static paradigms: (i) Fixed-size: The dominant paradigm relies on a uniform patch size across the sequence. This ranges from fine-grained point-wise tokenization [4, 31, 34] to coarser uni-size patching [29], which has been adopted by recent foundation models [40, 28]. While the latter improves upon the computational inefficiency of point-wise methods, both remain rigid. (ii) Multi-size: To capture broader temporal dynamics, models like Pathformer [8] and TTM [13] incorporate multiple predefined patch sizes. However, these approaches apply selected granularities uniformly across the entire sequence based on global features rather than local information. This rigidity forces the model to allocate equal computational cost to segments of varying complexity. This imbalance leads to suboptimal parameter efficiency and limited generalization. In contrast, Kairos employs a mixture-of-size strategy. By allocating fewer tokens to simple patterns and more to complex regions, Kairos enhances parameter efficiency while maintaining robust generalization.
@@
Position Embedding. Due to the position-unaware characteristic of self-attention mechanism, Transformers [36] rely on position embeddings (PEs) to model temporal information. However, most TSFMs adopt existing designs for PE in Natural Language Processing (NLP) domains [11, 34], which typically emphasize long-term decay through methods like sinusoidal calculations or by suppressing long-range positional information [36, 35]. By imposing a uniform temporal scale, these PEs struggle to model the heterogeneous temporal relationships of time series. Existing adaptive PE methods in NLP [45, 23] are also insufficient, as they target context extrapolation rather than the complex temporal structures of time series. Therefore, ElasTST [44] proposed a tunable RoPE to better adapt to time series data. However, its adaptation uses per-dataset training rather than dynamic adjustment. To address this limitation, we propose to dynamically modulate PE tailored to the intrinsic characteristics of each input time series, effectively decoupling temporal heterogeneity from model capacity, therefore enhancing the parameter efficiency and overall performance of TSFMs.
@@ 3.1 Overall Framework
We formulate our task as learning a time series forecasting function $f_{\phi}:\mathbb{R}^{T}\to\mathbb{R}^{H}$ parameterized by $\phi$. Given historical time series $\mathbf{X}\in\mathbb{R}^{T}$, time series forecasting is to predict the future horizon $\mathbf{Y}\in\mathbb{R}^{H}$. The learning objective is to minimize the discrepancy between the ground truth $\mathbf{Y}$ and the prediction $\hat{\mathbf{Y}}=f_{\phi}(\mathbf{X})$. We detail it in Appendix H.1.2.
@@ 3.2 Mixture-of-Size Encoder
Granularity. For a time series segment $\mathbf{x}$, we adopt a multi-granularity perspective. Granularity refers to the temporal resolution at which the segment is tokenized and encoded, as defined below.
@@
Step 1: Top-$K$ Granularity Routing. Applying encoding at an appropriate granularity level for each time series segment is non-trivial due to the inherent temporal heterogeneity of time series data. To address this challenge, we introduce a dynamic routing mechanism equipped with a lightweight router that adaptively selects an optimal subset of granularity levels for subsequent encoding of a given segment. Inspired by the Mixture-of-Experts (MoE) paradigm [19, 33], we instantiate $G$ granularity-specific experts, each corresponding to a distinct granularity level, together with $Z\in\{1,\dots,K-1\}$ computation-free null experts. The inclusion of null experts allows the model to implicitly skip encoding at certain granularities, enabling adaptive control over the effective number of active granularities for each input segment. For each expert $i\in\{1,\dots,G+Z\}$, the routing weights are computed via a learnable affinity score $s_{i}=\mathbf{w}_{i}^{\top}\mathbf{x}$, normalized with a bias-corrected softmax $\tilde{s}_{i}=\left(\exp(s_{i}+\boldsymbol{b}_{i})\right)/\left(\sum_{j=1}^{G+Z}\exp(s_{j}+\boldsymbol{b}_{j})\right),$ where $\mathbf{w}_{i}$ denotes learnable parameters and $\boldsymbol{b}_{i}$ is a bias for auxiliary-loss-free load balancing [24] (details in Appendix H.3.1). The router then activates the indices $\mathcal{I}_{\mathrm{valid}}=\operatorname*{argtop\text{-}K}_{i\in[G+Z]}\penalty 10000\ \tilde{s}_{i}\cap\{1,\dots,G\}$. By allowing null experts to dynamically prune the number of active experts, Kairos dynamically adapts the effective resolution to local information while maintaining a sparse, parameter-efficient architecture.
@@
where $\pi_{i}$ denotes a granularity-specific projection function, implemented as a two-layer MLP.
Step 3: Mixture-of-Size Fusion. In the final encoder stage, we fuse the embeddings produced by the activated granularity experts into a unified token sequence, serving as the input to the Heterogeneity-Aware Transformer. First, for a time-series segment $\mathbf{x}$, we align all granularity-specific embeddings $\boldsymbol{E}_{i}$ to the finest activated resolution $P_{\min}=\min\{P_{i}\mid i\in\mathcal{I}_{\mathrm{valid}}\},$ yielding a target sequence length $M=\lceil P/P_{\min}\rceil$. This alignment is crucial to eliminate representation bias arising from unequal sequence lengths induced by different patch sizes across granularities. To enable efficient alignment, we design the patch sizes to be nested, i.e., $P_{i+1}$ is divisible by $P_{i}$. Under this constraint, each embedding sequence $\boldsymbol{E}_{i}\in\mathbb{R}^{M_{i}\times D_{h}}$ can be deterministically expanded to the target resolution via repetition (details and an example in Appendix H.3.2): $\tilde{\boldsymbol{E}}_{i}=\mathrm{Expand}(\boldsymbol{E}_{i})\in\mathbb{R}^{M\times D_{h}}.$ The fused representation is obtained by a gated weighted summation over the activated granularities: $\bar{\boldsymbol{E}}=\sum_{i\in\mathcal{I}_{\mathrm{valid}}}c_{i}\,\tilde{\boldsymbol{E}}_{i},c_{i}=\tilde{s}_{i}/{\sum_{j\in\mathcal{I}_{\mathrm{valid}}}\tilde{s}_{j}}.$
@@
Unlike prior approaches such as Chen et al. [8], Zhang et al. [44], which impose a static multi-scale partitioning over the entire time series, our method (i) dynamically adapts the tokenization granularity at the segment level. (ii) decouples temporal heterogeneity from a fixed representational scale by explicitly and dynamically assigning patch size via routing and multi-granularity encoding, enabling more effective intra-patch representation learning. This adaptive behavior cannot be achieved via globally predefined scales applied statically and is empirically validated in Section 4.4.1 and Appendix E. As a result, varying temporal patterns are modeled via the lightweight Mixture-of-Size Encoder, reducing the burden of subsequent model parameterization.